# Dialectal Coverage And Generalization in Arabic Speech Recognition

### v2

@@ 4.3 Zero-Shot & Fine-Tuning Results
To quantify the effect of dialectal pre-training, we compare fine-tuning performance using the same data sets on both v1 and v2. We first fine-tune models on MSA data using the MGB2 benchmark dataset. We test the model performance on other dialects without dialectal fine-tuning (zero-shot) and with dialectal fine-tuning. The results are shown in Table 6. On average, we see improvement in performance in both zero-shot and fine-tuning experiments using v2 compared to v1. We also see that both v1 and v2 perform better than Whisper and MMS in zero-shot settings in most cases. There are some exceptions, such as in KUW, where Whisper performs better than all other models, including the fine-tuned models, but in all other cases v2 performs better. This underscores the advantage of monolingual models compared to multi-lingual performance, as observed in Toyin et al. (2023) and Radford et al. (2023). In addition, the results underscore the importance of dialectal coverage in pre-training: the cases where v2 does not perform better than v1 are all dialects for which pre-training data are limited, such as TUN (no pre-training data) and JOR (smallest dialect size in pre-training).
We conclude from this section that utilizing more diverse data during the unspervised pre-training stage is recommended for optimizing downstream tasks across more dialectal variants.
@@ 5 Joint Models & Dialect ID
We assess the relative effectiveness of individual dialectal fine-tuning compared with joint dialect fine-tuning. To that end, we joined multiple dialectal train sets as shown in Table 7.
From MASC, we excluded ALG, YEM, SUD for zero-shot evaluation. The resulting joint corpus consists of 12 dialects including MSA, with approximately 1,501 training hours in total. We train a single joint model using this data.
@@
Table 7: Datasets used to train the joint model.
Fine-Tuning Data zero-shot Fine-Tuning Joint Dialects
[…]
@@
Table 8: WER (%) of various models compared with joint dialectal fine-tuning with different dialect ID strategies
@@ 5.1 Dialect ID
For inference, we experimented with two strategies: (1) Transcribing with dialect forcing, where we manually add the dialect ID to condition the decoder output; the decoder is forced to start predictions with the tokens <S> DIALECT . (2) Transcribing with dialect inference, where we let the model predict the dialect token automatically. We use this approach for zero-shot ASR on unseen dialects.
The results of the models trained with joint dialects compared to models trained on MGB2 and QASR are shown in Table 8. Note that both MGB2 and QASR contain mostly MSA, but also a small amounts of various dialects, but their exact distribution is unknown. We also show the fine-tuning results from Table 6 for easy comparison. We see that joint modeling results in improvement for low-resource dialects, including: JOR, TUN, and KUW, but degrades performance of the high-resource SYR and EGY dialects. Interestingly, dialect forcing, results in average performance similar to target fine-tuning, and dialect inference resulted in the best performance overall, followed by the model without dialect ID. We surmise that the model learns dialectal patterns that do not perfectly align with the dialect ID as indicated in the training data. Since the dialect IDs are coarse country-level approximations, letting the model infer the dialect that best aligns with the speech is the best approach for most cases. Many dialectal sets, such as SYR and SAU, contain a lot of MSA utterances that are incorrectly identified as dialectal, and low-resource dialects, such as KUW, are predicted as their closest high-resource variant, such as SAU.
Figure 3 illustrate dialect inference errors. Note that the number of errors is proportional to the test data size. The overall dialect identification performance is around 90%.
@@ 6 Effect of Data Adaptation
dialects Fine-Tuning Set
@@
ALG SUD YEM
WER (%) WER (%) WER (%)
[…]
v2 Joint 45.20 40.69 33.08
v2 joint (w. dialect inference) 47.12 40.15 31.84
Table 11: WER of various models on held-out dialects.
In the above experiments, we followed the strategy of initializing the models by first fine-tuning on MSA data. In most cases, we used MGB2 as the base model, following previously established results on Egyptian ASR Ali et al. (2017). This adaptation approach is meant to enhance the performance on low-resource dialects, facilitating faster convergence with limited training samples. However, as pre-training on more diverse sets proved to be effective, adaptation on more diverse data is also likely to be fruitful. As observed in Table 8, models trained on QASR resulted in far better zero-shot performance, approaching the performance of joint-dialects models. This is attributed to the fact that QASR is both larger in size and known to have more dialectal data compared to MGB2 (but both have no documented statistics of dialectal coverage). To validate this observation, we experimented with dialectal fine-tuning adapted from two variants: one based on MGB2 and one based on QASR Mubarak et al. (2021), followed by dialect-specific fine-tuning. Table 5 shows fine-tuning results with no adaptation (directly fine-tuning on the target dialect), compared with starting from MGB2 or QASR. First, our results corroborate the previous findings that adapting models from MSA results in large reduction in error rates. In all except the Mixat dataset, starting from QASR results in better performance compared to MGB2. However, the difference is negligible except on MGB3 Egyptian set (around 10% absolute WER reduction). There are two factors that we speculate underline this result: The small size of the MGB3 set, and the existence of Egyptian dialect in the QASR corpus more substantially than the other dialects. Overall, using the QASR dataset as a basis for adapting dialectal models is recommended as it improves or maintains performance.
7 Cross-Dialectal Performance
@@
Table 10 shows the cross-dialectal performance, where models trained on a single target dialect are tested on other dialects, including the three held-out sets: ALG, SUD, and YEM. We also show the performance of zero-shot and joint models on the held-out dialects in Table 11. In most cases, the best performance is achieved by the model trained on the same target dialect (the diagonal in Table 10). However, for low-resource dialects, like KUW and PAL, the model trained on SAU achieved the lowest WER. This is likely a result of the large size of the SAU train set and the wide geographical area and dialectal variants it covers. The joint models perform best on the held-out dialects. Overall, the performance on the held-out sets is on a par with the low-resource dialects, with WER above 30%. Curiously, all models perform well on the SYR test set; upon close inspection, we found that the set consists mostly of MSA utterances, which explains the result since all models are adapted from MSA.
8 Conclusions
We presented the largest study on dialectal Arabic ASR to empirically demonstrate the effect of various training paradigms on ASR performance. We compared models pre-trained with and without dialects, in high, low, and medium-resource settings, in addition to zero-shot. We find that overall, dialectal pre-training improves performance in zero-shot and low-resource cases, and mostly maintains performance on MSA and high-resource dialects. We also find that all dialects benefit from adaptation of models pre-fine-tuned on MSA, and this effect is most noticeable for low and medium-resource dialects. Adaptation from larger and more varied datasets, like QASR, results in improved performance, particularly in ’zero-shot’ settings. The latter is likely influenced by the presence of dialectal data in QASR, so it is not pure zero-shot as the exact composition of the dataset is not documented. We experimented with more controlled multi-dialectal fine-tuning, where we joined the train sets of 12 dialects. In this case, we observe additional improvements compared to QASR in spite of having smaller size overall, where we achieve on average at least the same performance as the target-dialect fine-tuning setting, and the best performance on held-out dialects. Interestingly, while using dialect ID in decoding is effective, forcing the dialect ID results in worse performance compared to dialect inference. Finally, while joint training results in improved performance for the medium and low-resource dialects, target-dialect fine-tuning is more effective for high-resource dialects. In light of the above experiments and results, we recommend the following: (1) include as many dialectal variants as possible in the pre-training stage, (2) fine-tune ASR on large sets of joint dialects, preferably with the dialect ID inference, then (3) fine-tune separate models for high-resource dialects with more than 100 hours of labeled data. To enable easier adoption and further experiments, we will release the pre-trained v2 checkpoint, the fine-tuned MGB2 and QASR models, and the joint model with dialect tokens.
Limitations
One of the limitations in dialect-related work is the coarse classification of dialect IDs; dialects in our datasets are classified by regions or countries, whereas actual dialectal variations are far more fine-grained. For example, the Saudi dataset, SADA, covers a large geographical area and many dialects, but it is considered as one dialect based on our classification. Moreover, the way the datasets are collected do not guarantee that the data are indeed dialectal. For instance, with manual inspection of the Syrian test and dev sets from MASC, we observed that all instances are in MSA rather than Syrian dialects. In addition, Arabic dialects are spoken varieties that do not have a standard spelling system. This results in large variations in transcriptions, but standard WER does not account for these variations, resulting in more pessimistic results. With the exception of the MGB3 and MGB5 benchmarks where we report average and multi-reference WER across 4 references, all datasets have only a single reference. Code-switching is also observed in some datasets, such as Mixat, but the model has limited ability to recognize multi-lingual code-switching as it is only trained on Arabic. Given small amount of English code-switching in the pre-training data, the model does have limited code-switching recognition, but it can be improved by intentionally incorporating code-switching datasets for pre-training and fine-tuning. The combination of multi-lingual and multi-dialectal training has not been explored in this work.
Appendix A Inference examples
Figure 5: Examples from the Mixat & TARIC-SLU dataset illustrating the zero-shot performance of v1 and v2 fine-tuned on MGB2 vs. v2 fine-tuned on QASR.
Figure 6: Example of ASR outputs on held-out dialects from dialect-specific fine-tuning.
Table 4 lists examples of ASR outputs using the dialect-specific fine-tuned models. Note that the ‘errors’ in SAU, EGY, and JOR examples are in fact alternative spellings. Figure 5 compares zero-shot v1 and v2 (adapted from MGB2) and v2 (adapted from QASR) performance from the Mixat (UAE) and TARIC (Tunisia) datasets.

### v3

@@ 4.3 Zero-Shot & Fine-Tuning Results
To further quantify the effect of dialectal pre-training, we evaluate the performance of our model across different datasets. We first fine-tune models on MSA using MGB2 dataset. We test the model performance on dialects directly (zero-shot) and with dialectal fine-tuning. The results are shown in Table 5. On average, we see improvements in performance in both zero-shot and fine-tuning experiments using dialectal pretraining (v2) compared to MSA-centric pretraining (v1). We also see that both models perform better than Whisper and MMS in zero-shot settings in most cases. There are some exceptions, such as in KUW, where Whisper performs better than all other models, including the fine-tuned models, but in most cases v2 performs best. This underscores the advantage of monolingual models compared to multilingual performance, as observed in Toyin et al. (2023) and Radford et al. (2023). In addition, the results underscore the importance of dialectal coverage in pretraining: the cases where v2 performs worse than v1 are all dialects for which pretraining data are limited, such as TUN (no pretraining data) and JOR (smallest dialect size in pretraining).
@@ 5 Joint Models & Dialect ID
So far, models were first fine-tuned on MSA, followed by additional fine-tuning on each target dialect. This results in a separate model per dialect, which incurs memory costs and may have practical limitations as it requires advance knowledge of the dialect ID for deploying the correct model.
In this section, we assess the relative effectiveness of individual dialectal fine-tuning compared with joint dialect fine-tuning, where we train a single model for all dialects. To that end, we joined multiple dialectal train sets from MASC, as shown in Table 6. We excluded ALG, YEM, SUD for zero-shot evaluation. The resulting joint corpus consists of 12 dialects including MSA, with approximately 1,501 hours in total. We fine-tuned a single joint model using this data.
@@
Table 6: Datasets used to train the joint models.
Approach Zero-Shot Fine-Tuning No Dialect ID Dialect Forcing Dialect Inference
Fine-Tuning Data MGB2 QASR MGB2$\rightarrow$ Target Joint Mutli-Dialectal Set (Table 6)
@@
Table 7: WER (%) of various models compared with joint dialectal fine-tuning with different dialect ID strategies.
@@ 5.1 Dialect ID
For inference, we experimented with two strategies: (1) Transcribing with dialect forcing, where we manually add the dialect ID to condition the decoder output; the decoder is forced to start predictions with the tokens <S> DIALECT . (2) Transcribing with dialect inference, where we let the model predict the dialect token. We use this approach for zero-shot ASR on unseen dialects (Table 9).
The results of the models trained with joint dialects compared to models trained on MGB2 and QASR are shown in Table 7. Note that both MGB2 and QASR contain mostly MSA, but also a small amounts of various dialects, but their exact distribution is unknown. We also reproduce the fine-tuning results from Table 5 for easy comparison. We see that joint modeling results in improvement for low-resource dialects, including: JOR, TUN, and KUW, but degrades performance of the high-resource SYR and EGY dialects. Interestingly, dialect forcing was worse on average than joint modeling with no dialect ID, while dialect inference resulted in the best performance overall. We surmise that the model learns dialectal patterns that do not perfectly align with the dialect ID as indicated in the training data. Since the dialect IDs are coarse country-level approximations, letting the model infer the dialect based on the speech is the best approach for most cases. Many dialectal sets, such as SYR and SAU, contain a lot of MSA utterances that are incorrectly identified as dialectal.
Figure 3 illustrates dialect inference errors. Note that the number of errors is proportional to the test data size. The overall dialect identification performance is around 90%. Some low-resource dialects, such as KUW, are predicted as their closest high-resource variant, such as SAU, resulting in worse performance compared to joint models without a dialect ID, but on average, dialect forcing leads to the lowest WER.
@@ 6 Effect of Data Adaptation
In the above experiments, we followed the strategy of initializing the models by first fine-tuning on MSA data. In most cases, we used MGB2 as the base model, following previously established results on Egyptian ASR Ali et al. (2017). This adaptation approach is meant to enhance the performance on low-resource dialects, facilitating faster convergence with limited training samples. However, as pre-training on more diverse sets proved to be effective, adaptation on more diverse data is also likely to be fruitful. As observed in Table 7, models trained on QASR resulted in far better zero-shot performance, approaching the performance of joint-dialects models. This is attributed to the fact that QASR is both larger in size and known to have more dialectal data compared to MGB2 (but both have no documented statistics of dialectal coverage). To validate this observation, we experimented with dialectal fine-tuning adapted from two variants: one based on MGB2 and one based on QASR Mubarak et al. (2021), followed by dialect-specific fine-tuning. Table 4 shows fine-tuning results with no adaptation (directly fine-tuning on the target dialect), compared with starting from MGB2 or QASR. First, our results corroborate the previous findings that adapting models from MSA results in large reduction in error rates. In all except the Mixat dataset, starting from QASR results in better performance compared to MGB2. However, the difference is negligible except on MGB3 Egyptian set (around 6% absolute WER reduction). There are two factors that we speculate underline this result: The small size of the MGB3 set, and the existence of Egyptian dialect in the QASR corpus more substantially than the other dialects. Overall, using the QASR dataset as a basis for adapting dialectal models is recommended as it improves or maintains performance.
7 Zero-Shot Performance
System Dialect ALG SUD YEM
[…]
Table 9: WERof various models on held-out dialects.
We show the zero-shot performance on the three held-out sets: ALG, SUD, and YEM. We compare the baseline, v1, with multi-dialectal pretraining (v2). We also compare models fine-tuned on MGB2, QASR, or our joint dialectal set. The results are shown in Table 9. Our model achieves slightly lower error rates compared to v1, even when fine-tuned on the same MGB2 set. Better performance is achieved with QASR, which includes some dialectal data. The joint dialectal fine-tuning achieves the best performance on the held-out dialects. In general, performance in held-out sets is on a par with low-resource dialects, with WER above 30%. Table 13 in the Appendix shows the zero-shot performance after fine-tuning on a single target dialect.
8 Code-Switching Performance
The models analyzed so far were trained exclusively on Arabic data. While small amounts of code-switching (CS) exist in these sources, they are insufficient to learn the characteristics of the embedded languages. Large multi-lingual models are generally more effective on CS data Kadaoui et al. (2024), even if they are less competent on monolingual Arabic. To make our models more inclusive, improving performance in the presence of code-switching is necessary. To that effect, we train a multilingual version of the model (we will refer to this as v3). The pre-training data for this version are listed in Table 12 in the Appendix. We test v3 against v1 and v2 on available CS data for Arabic: ArZN Al-Sabbagh (2024) for Egyptian-English speech, Mixat Al Ali and Aldarmaki (2024) for Emirati-English speech, and TunSwich Abdallah et al. (2024) for Tunisian-French speech. We also train a joint multi-lingual model (without dialect or language ID). In addition to the datasets described in Table 6, we add the multi-lingual and code-switching data shown in Table 10.
In Table 11, we show the performance of models finetuned directly from the pretrained checkpoints, or finetuned from existing ASR checkpoints (MGB2 checkpoint for v1, joint multi-dialectal checkpoint for v2, and joint multi-lingual checkpoint for v3). First, for models fine-tuned directly on the target set, we observe that multilingual pretraining significantly improves performance across all CS test sets, resulting in more than 10% absolute reductions in WER for all test sets. This clearly illustrates the advantage of multi-lingual pretraining in code-switching scenarios. We also evaluated models initialized from the joint models followed by target fine-tuning on the CS train sets, and this reduced error rates further. The best performing model is the joint multilingual v3 mode, with 4 to 7% absolute reduction in WER compared to the second best model. We show examples of ASR outputs from the three CS datasets using the various models in Figure 6 in the Appendix.
Languages Hours Words Source
EN 1601.92 hrs 10.35 M CommonVoice
FR 732.02 hrs 5.03 M CommonVoice
SP 408.34 hrs 2.79 M CommonVoice
TUN-FR 10.89 hrs 70.86 K TunSwitch
UAE-EN 8.97 hrs 57.82 K Mixat
EGY-EN 5.61 hrs 52.00 K ArzEn
Table 10: Additional datasets used to train the joint multilingual model.
Adaptation data - - - MSA Dialectal Multilingual
Test Set v1 v2 v3 v1 v2 v3
MGB2 13.42 12.5 13.0 - - -
ArzEn 43.21 77.59 35.26 34.85 33.71 27.43
TunSwitch 53.85 101.94 40.68 43.87 43.59 36.66
Mixat 42.50 92.41 34.27 27.07 25.73 21.66
Table 11: ASR Results using the various checkpoints: v1, v2 and v3. We compare models trained directly from the pretrained checkpoint vs. starting with an ASR checkpoint trained with the specified adaptation data: MSA adaptation using the MGB2 dataset; Dialectal adaptation using data listed in Table 6; Multilingual adaptation using data from Table 6 and Table 10.
Language Interference:
We test the effect of multi-lingual pre-training on MSA performance. Language interference is known to negatively affect monolingual performance Toyin et al. (2023), so we test our multilingual model on the MGB2 benchmark to quantify this effect (see Table 11). The model achieves 13.0% WER, which is indeed worse than the SOTA result we achieve with the Arabic-only model (see Table 2), but the difference at 0.5% absolute WER is rather small. When it comes to dialects, however, we find that language interference has a significant negative effect, resulting in 4% to 16% absolute increase in error rates, as shown in Figure 4.
Figure 4: WER (%) and absolute difference on a subset of dialects, comparing our joint dialectal fine-tuning vs. joint multi-lingual fine-tuning on Arabic dialects.
9 Conclusions
We presented the largest study on dialectal Arabic ASR to empirically demonstrate the effect of various training paradigms on ASR performance. We compared models pre-trained with and without dialects, in high, low, and medium-resource settings, in addition to zero-shot. We find that overall, dialectal pre-training improves performance in zero-shot and low-resource cases, and mostly maintains performance on MSA and high-resource dialects. We also find that all dialects benefit from adaptation of models pre-fine-tuned on MSA, and this effect is most noticeable for low and medium-resource dialects. We experimented with multi-dialectal fine-tuning, where we joined the train sets of 12 dialects. We observe performance improvements on average, and at least the same performance as the target-dialect fine-tuning setting, and the best performance on held-out dialects. Interestingly, while using dialect ID in decoding is effective, forcing the dialect ID results in worse performance compared to dialect inference. While joint training results in improved performance for the medium and low-resource dialects, target-dialect fine-tuning is more effective for high-resource dialects. Finally, we experimented with multi-lingual pre-training and fine-tuning for improving performance on code-switched utterances, and achieved significant reductions in error rates on all available test sets. However, reductions in monolingual performance were also observed due to language interference, particularly for dialects, were WER increased by 4% to 16% for some dialects. To enable easier adoption and further experiments, we released the pretrained dialectal and multilingual checkpoints, the fine-tuned MGB2 models, and the joint dialectal and multilingual ASR models.
Limitations
One of the limitations in dialect-related work is the coarse classification of dialect IDs; dialects in our datasets are classified by regions or countries, whereas actual dialectal variations are far more fine-grained. For example, the Saudi dataset, SADA, covers a large geographical area and many dialects, but it is considered as one dialect based on our classification. Moreover, the way the datasets are collected do not guarantee that the data are indeed dialectal. For instance, with manual inspection of the Syrian test and dev sets from MASC, we observed that all instances are in MSA rather than Syrian dialects. In addition, Arabic dialects are spoken varieties that do not have a standard spelling system. This results in large variations in transcriptions, but standard WER does not account for these variations, resulting in more pessimistic results. With the exception of the MGB3 and MGB5 benchmarks where we report average and multi-reference WER across 4 references, all datasets have only a single reference.
Appendix A Pre-Training Dataset Statistics
Table 12 shows the complete list of datasets used for pre-training v1, v2, and v3.
Dataset Dialect Hours Words v1 v2 v3
ASC MSA 3.7 hrs 20.58 K $\checkmark$ $\checkmark$
ArzEn${}^{\texttt{[cs]}}$ EGY 5.61 hrs 52.00 K $\checkmark$
CommonVoice Dialect Mix 133.24 hrs 494.83 K $\checkmark$ $\checkmark$
CommonVoice ENG 1601.92 10.35 M $\checkmark$
FR 732.02 5.03 M $\checkmark$
ES 408.34 2.79 M $\checkmark$
ClArTTS CA 12 hrs 76.31 K $\checkmark$ $\checkmark$
MASC EGY 175.36 hrs 1.03 M $\checkmark$
IRA 17.37 hrs 121.12 K $\checkmark$
JOR 42.21 hrs 330.83 K $\checkmark$
KUW 4.04 hrs 32.37 K $\checkmark$
LEB 25.20 hrs 155.76 K $\checkmark$
MOR 8.60 hrs 58.38 K $\checkmark$
MSA 612.28 hrs 3.80 M $\checkmark$
PAL 6.17 hrs 45.35 K $\checkmark$
SAU 452.24 hrs 301.92 K $\checkmark$
SYR 211.33 hrs 1.06 M $\checkmark$
TUN 12.17 hrs 34.34 K $\checkmark$
UAE 9.87 hrs 6.42 K $\checkmark$
MGB2 Mostly MSA 1000 hrs 7.31 M $\checkmark$ $\checkmark$ $\checkmark$
QASR Mostly MSA 2000 hrs 13.33 M $\checkmark$
MGB3[ASR] EGY 2.83 hrs 18.93 K $\checkmark$
MGB3[ADI] EGY 11.15 hrs $\checkmark$ $\checkmark$
GLF 8.92 hrs $\checkmark$ $\checkmark$
LAV 9.27 hrs $\checkmark$ $\checkmark$
MSA 9.39 hrs $\checkmark$ $\checkmark$
NOR 9.49 hrs $\checkmark$ $\checkmark$
MGB5[ASR] MOR 115.7hrs $\checkmark$
MGB5[ADI] ALG 115.7hrs $\checkmark$ $\checkmark$
EGY 451.1 hrs $\checkmark$ $\checkmark$
IRA 815.8 hrs $\checkmark$ $\checkmark$
JOR 25.9 hrs $\checkmark$ $\checkmark$
KSA 186.1 hrs $\checkmark$ $\checkmark$
KUW 108.2 hrs $\checkmark$ $\checkmark$
LEB 116.8 hrs $\checkmark$ $\checkmark$
LIB 127.4 hrs $\checkmark$ $\checkmark$
MAU 456.4 hrs $\checkmark$ $\checkmark$
MOR 57.8 hrs $\checkmark$ $\checkmark$
OMA 58.5 hrs $\checkmark$ $\checkmark$
PAL 121.4 hrs $\checkmark$ $\checkmark$
QAT 62.3 hrs $\checkmark$ $\checkmark$
SUD 47.7 hrs $\checkmark$ $\checkmark$
SYR 119.5 hrs $\checkmark$ $\checkmark$
UAE 108.4 hrs $\checkmark$ $\checkmark$
YEM 53.4 hrs $\checkmark$ $\checkmark$
Mixat${}^{\texttt{[cs]}}$ UAE 9.97 hrs 57.82 K $\checkmark$
ParallelCorp EGY 32 hrs 48.56 K $\checkmark$
GLF 32 hrs 27.26 K $\checkmark$
LEV 32 hrs 18.43 K $\checkmark$
MSA 32 hrs 30.66 K $\checkmark$
MADAR MOR ALG TUN 532.37K $\checkmark$ $\checkmark$
LIB EGY LEV
IRA GLF YEM
NADI ALG BAH EGY 702.67K $\checkmark$ $\checkmark$
IRA JOR KUW
LEB LIB MOR
OMN PAL QAT
SAU SUD SYR
TUN UAE YEM
SADA SAU 417.63 hrs 3.26M $\checkmark$
TARIC-SLU TUN 6.81 hrs 53.48K $\checkmark$
TunSwitch${}^{\texttt{[cs]}}$ TUN 10.89 hrs 70.86K $\checkmark$
VoxBlink Dialect Mix 19.92 hrs $\checkmark$
Table 12: Summary of Dataset Statistics for Pre-Training: Hours of Audio, Word Counts, and Associated Dialects. *${}^{\texttt{[cs]}}$ refers to Code Switching datasets. *${}^{\texttt{[txt]}}$ refers to textual datasets.
Appendix B Inference examples
Figure 5 lists examples of ASR outputs using the dialect-specific fine-tuned models. Note that the ‘errors’ in SAU, EGY, and JOR examples are in fact alternative spellings.
Figure 6: Examples of code-switching samples from each of the code-switching datasest using (a) pre-training checkpoints fine-tuned on CS data directly, and (2) adapted models after fine-tuning on general data: v1 (MGB2), v2 (Dialectal data from Table 6), and v3 (dialectal data from Table 6 + Multilingual and CS data from Table 10).
Dialect Fine-Tuning Set
@@ Appendix B Inference examples
Figure 6 shows example outputs from each model on the code-switching datasets: ArzEn (Egyptian-English), TunSwitch (Tunisian-French), and Mixat (Emirati-English).
Appendix C Cross-Dialectal Performance
@@
Table 13 shows the cross-dialectal performance, where models trained on a single target dialect are tested on other dialects, including the three held-out sets: ALG, SUD, and YEM. In most cases, the best performance is achieved by the model trained on the same target dialect (the diagonal in Table 13). However, for low-resource dialects, like KUW and PAL, the model trained on SAU achieved the lowest WER. This is likely a result of the large size of the SAU train set and the wide geographical area and dialectal variants it covers. Curiously, all models perform well on the SYR test set; upon close inspection, we found that the set consists mostly of MSA utterances, which explains the result since all models are adapted from MSA.